Today, we learned that brachydactyly is a genetically dominant trait that caused malformation of the hands. However, nobody in the class actually had brachydactyly! The most common form is brachydactyly A3 (BDA3), and is typically seen in only 2% of people, so it is not surprising that nobody in this class (29 students) would have the trait. What are the expected genotype frequencies IF humans are at Hardy-Weinberg equilibrium for BDA3? I have never had a student with brachydactyl: that’s over 1000 people. How many students should I have had with this trait? Does this imply that the human population is evolving with regard to brachydactyly, and if so, which form of selection should this represent? WHY might this be occurring (how does BDA3 affect fitness)?
Answer #1

Genotype frequency of brachydactyly trait = 0.02 = p2+2pq (Note that all heterozygous individuals will have brachydactyly because brachydyly is a dominant trait.)

p2 + 2pq +q2 =1

q2 = 1 - 0.02= 0.98 = Genotype frequency of person not having brachydactyly.

Out of 1000, it could have been possible to have the trait in 20 people. But since 98% people are autosomal recessive, so they will mate with persons who are autosomal recessive. I mean to say that heterozygous persons are not present in 98% of the population. So, the gene for brachydactyly becomes extremely rare.

This type of selection is known as stabilising.